For each molecule, predict the splitting pattern or patterns that would be seen in the peaks of
the high-resolution proton NMR spectrum.
a CH 3 CH 2 Cl

(1 mark)

b CH 3 COCH 3

(1 mark)

c CH 3 CHO

(1 mark)

d CH 3 CHClCH 3
1
Expert's answer
2020-10-09T13:46:34-0400

We know that in proton NMR spectroscopy the number of splitting for a particular type of chemically equivalent proton is,

Number of splitting = (2NiI + 1) when only one type of proton is present

Ni = number of neighborhood proton

I = nuclear spin quantum number

= (1/2)

So, (2NiI + 1) = (Ni+1)

Ni = number of neighborhood proton

I = nuclear spin quantum number

= (1/2)

So, (2NiI + 1) = (Ni+1)


When more than one type of protons are present, (2NiI + 1)×(2NjI + 1)×....

Ni or Nj = number of neighborhood proton

I = nuclear spin quantum number

I = 1/2

So, (2NiI + 1)×(2NjI + 1)×....

= (Ni+ 1)×(Nj+ 1)×....




a. CH3 CH2Cl

CH3

Ni= 2.

Splitting = (2+1)

=3

A triplet


CH2

Ni= 3

Splitting = (3+1)

= 4

A quartet





b. CH2 COCH3


CH3

Ni= 0

Splitting = (0+1)

=1

No splitting


COCH3

Ni= 0

Splitting = (0+1)

= 1

No splitting




c. CH3 CHO


CH3

Ni= 1

Splitting = (1+1)

=2

A doublet


CHO

Ni= 3

Splitting = (3+1)

= 4

A quartet




d. CH3 CHCl CH3


CH3

Ni= 1

Splitting

(1+1) =2

A doublet

CHCl

Ni= 3, Nj =3.

Splitting

(3+1)×(3+1)= 16

A multiplate

CH3

Ni =1

Splitting

=(1+1)

= 2

A doublet